# OCR Hinders RAG: Evaluating the Cascading Impact of OCR on Retrieval-Augmented Generation

### v1

@@ Abstract
Retrieval-augmented Generation (RAG) enhances Large Language Models (LLMs) by integrating external knowledge to reduce hallucinations and incorporate up-to-date information without retraining. As an essential part of RAG, external knowledge bases are commonly built by extracting structured data from unstructured PDF documents using Optical Character Recognition (OCR). However, given the imperfect prediction of OCR and the inherent non-uniform representation of structured data, knowledge bases inevitably contain various OCR noises. In this paper, we introduce OHRBench, the first benchmark for understanding the cascading impact of OCR on RAG systems. OHRBench includes 350 carefully selected unstructured PDF documents from six real-world RAG application domains, along with Q&As derived from multimodal elements in documents, challenging existing OCR solutions used for RAG. To better understand OCR’s impact on RAG systems, we identify two primary types of OCR noise: Semantic Noise and Formatting Noise and apply perturbation to generate a set of structured data with varying degrees of each OCR noise. Using OHRBench, we first conduct a comprehensive evaluation of current OCR solutions and reveal that none is competent for constructing high-quality knowledge bases for RAG systems. We then systematically evaluate the impact of these two noise types and demonstrate the vulnerability of RAG systems. Furthermore, we discuss the potential of employing Vision-Language Models (VLMs) without OCR in RAG systems. Code: https://github.com/opendatalab/OHR-Bench
footnotetext: These authors contributed equally to this work.footnotetext: Project leader.footnotetext: Corresponding author (heconghui@pjlab.org.cn).
@@ 1 Introduction
As an essential component of RAG systems, the knowledge base defines the scope and quality of documents that RAG can access. Given that a vast amount of real-world knowledge resides in unstructured documents, such as scanned PDFs, constructing an external knowledge base often relies on Optical Character Recognition (OCR) to parse structured data from these unstructured PDF documents [39, 14]. For instance, MinerU [32] takes raw PDFs as input and extracts plain text, formulas, and tables into structured formats for subsequent RAG applications. However, imperfect predictions of OCR and non-uniform representations of parsing results impair the construction of a high-quality knowledge base for RAG. To be specific, despite advancements in OCR [32, 34, 2], even the leading model cannot achieve perfect accuracy across all scenarios [22, 39]. Furthermore, structural data can inherently be parsed in non-uniform representation, such as Markdown or LaTeX. These issues introduce OCR noise in parsing results and diminish the quality of the knowledge base. Considering RAG is sensitive to input noise [10, 36, 7], recent works race on downstream RAG components, including more precise retrievers [4, 24, 19] and more advanced LLMs [1, 8, 37, 10]. However, the quality of OCR-based external knowledge bases and its cascading impact on these downstream RAG components have received less attention, which highlights a critical but unaddressed gap: the absence of benchmarks to assess OCR’s cascading impact on each component and entire system of RAG.
@@
Existing benchmarks either evaluate RAG holistically without fine-grained assessment [40], or consider limited OCR solutions without consideration of noise they introduce [14, 11]. To fill this gap, we introduce OHRBench, a question-answering benchmark designed to evaluate OCR’s cascading impact on each component and entire systems of RAG in two ways. First, OHRBench constructs a document-based RAG Q&A dataset. It encompasses complex unstructured PDF documents from six RAG real-world application areas, including Textbook, Law, Finance, Newspaper, Manual, and Academia. As shown in Tab. 1 and Fig. 2, more than 1800 pages of these PDFs contain multimodal elements, such as tables and formulas, or feature complex layouts, posing challenges for constructing high-quality knowledge bases for RAG systems. For questions, OHRBench provides a diverse set of Q&A pairs, with evidence sourced from diverse varied document elements, making it an ideal testbed for assessing the OCR’s impact on RAG performance. Second, OHRBench identifies two primary OCR noise types: Semantic Noise, resulting from prediction errors, and Formatting Noise, arising from non-uniform document element representation. By deriving samples from current OCR results and systematically perturbing ground truth structured data based on these samples, OHRBench generates a set of perturbed structured data with varying degrees of Semantic Noise and Formatting Noise that occur in OCR results, enabling exploration of the quantitative relationship between OCR noise and RAG performance.
@@
With OHRBench, we first conduct a comprehensive benchmark on current OCR solutions, including pipeline-based OCR systems [32, 25], end-to-end OCR models [2, 34] and Vision-Language Models for OCR [33, 5]. We reveal that even the best OCR solutions exhibit a performance gap of 7.5% at least, compared to the ground truth structured data, facilitating the importance of mitigating OCR noise in RAG systems. Further experiments on different types of OCR noise uncover that Semantic Noise consistently exert significant impact, while Formatting Noise affects specific retrievers and LLMs differently. offering valuable insights for developing RAG-tailored OCR solutions and noise robust models. Moreover, we also discuss employing Vision-Language Models (VLMs) in the RAG generation stage without OCR. We find that inputting image only cannot reach the performance of using OCR text, but by simply combining these two inputs, VLM can improve the performance by up to 24.5% and approach the performance of the ground truth text baseline, indicating its promising potential of applying VLMs in RAG systems.
@@
  • We present OHRBench, a question-answering benchmark designed to evaluate the impact of OCR on RAG systems. OHRBench includes various unstructured PDF documents from six RAG application areas and Q&As sourced from multimodal document elements, posing challenges to the employment of current OCR solutions in RAG systems.
  • We introduce two primary types of OCR noise in OHRBench: Semantic Noise and Formatting Noise. We offer perturbed structured data with varying levels of each noise type, facilitating further exploration of RAG-tailored OCR solutions on RAG systems.
  • We conduct a comprehensive evaluation of current OCR solutions and reveal that none of their extracted structured data is competent for constructing high-quality knowledge bases for RAG systems. Moreover, we provide a systematic analysis of the fine-grained and cascading impact of these two OCR noises on RAG performance, facilitating future development of RAG-tailored OCR solutions.
Figure 1: Construction of OHRBench and evaluation protocol. (1) Benchmark Dataset: we collect PDF documents from six domains, extract human-verified ground truth structured data, and generate Q&As derived from multimodal document elements. (2) RAG Knowledge Base: OCR Processed Structured Data for benchmarking current OCR solutions and Perturbed Structured Data for assessing the impact of different OCR noise types. (3) Evaluation of OCR impact on each component and the overall RAG system.
@@ 2.1 Retrieval-Augmented Generation
Retrieval-Augmented Generation (RAG) [18, 17, 27] integrates external knowledge into large language models (LLMs) to mitigate hallucinations. Although RAG technology enhances the generation capabilities of LLMs, it is notably sensitive to input noise. InfoRAG [36] characterizes this noise in RAG as incorrect and irrelevant content within retrieved text and reveals its impact on RAG performance. RAAT [10] further expands noise into relevant noise, counterfactual noise, and irrelevant types. However, these studies focus solely on chunk-level noise introduced during the retrieval stage and its effect on generation capabilities of LLMs, leaving the impact of noise derived from OCR results unexplored. GARAG [7] examines typographical errors, a form of OCR noise, but its scope is limited to plain text using only synthetic data, overlooking the variety of OCR noise encountered in real-world RAG applications. In this paper, we reveal the impact of noise introduced during the OCR stage, offering a comprehensive analysis of its impact on RAG systems.
@@ 2.2 Document parsing with OCR
OCR-based document parsing is a promising solution for structured data extraction from unstructured documents, facilitating applications like RAG. Current OCR solutions can be summarized into three categories, pipeline-based systems [32, 25], end-to-end models [34, 2, 21] and employing VLMs for OCR [5, 33, 15] Pipeline-based systems decompose OCR into multiple subtasks such as layout detection, text, formula, and table recognition, enabling fine-grained data extraction. End-to-end models take document images as input and output the overall recognition result in an end-to-end manner. Due to the achievement of VLMs on visual understanding, recent works have explored its application in OCR [22]. In this paper, we evaluate these OCR paradigms, examining their suitability for RAG applications across diverse, real-world document domains.
@@ 2.3 Benchmark and evaluation of Retrieval-Augmented Generation
Frameworks like RAGAS [9] and ARES [29] propose evaluating RAG systems based on context relevance, answer faithfulness, and answer relevance, using LLMs or fine-tuned discriminators for measurement. RGB [3] assesses the noise robustness, negative rejection, information integration, and counterfactual robustness of RAG in news data. MultiHop-RAG [30] focuses on multi-hop reasoning capabilities, while ClashEval [35] explores the context preference in conflicting evidence scenarios. However, these evaluations target specific components of RAG systems, and none of them discusses the impact of external knowledge base construction on RAG systems. Although UDA [14], VisRAG [38] and M3DocRAG [6] explore RAG’s effectiveness in document analysis, they consider limited OCR solutions and lack analysis of different OCR noise types. In this paper, we introduce OHRBench to comprehensively investigate OCR noise’s impact on RAG systems.
@@
Documents 350
- Domain $6$
- Total Pages $4012$
- Avg.Tokens $482.61/\text{page}$
- Avg.Data Type $1.45/\text{page}$
Questions $4598$
Avg.Question Token $18.56$
Avg.Answer Token $7.91$
@@
- Text $3334\ (72.5\%)$
- Formula $872\ (19.0\%)$
- Table $392\ (8.5\%)$
@@
- Short Text Answer $2482\ (54.4\%)$
- Judgment $849\ (18.5\%)$
- Numeric $894\ (19.4\%)$
- Formula $354\ (7.7\%)$
@@
Figure 2: The layout of documents in OHRBench is complex. Each number indicates the count of PDF pages with that attribute. Criteria for these attributes can be found in Appendix Sec. II
@@ 3 OHRBench
Our OHRBench consists of 1) a number of unstructured PDF documents from six real-world RAG applications and Q&A pairs derived from multimodal elements in these documents, and 2) perturbed structured data based on ground truth with varying degrees of OCR noises.  Fig. 1 illustrates the construction of OHRBench. We will now delve into the details of each component.
@@ 3.1 Data collection
According to [39, 20], extracting structured data from multimodal document elements like formulas and tables poses significant challenges to current OCR solutions. Considering the practical application scenarios of RAG and the challenging field of OCR, we compile a PDF document collection representing six common RAG application scenarios: Textbook, Law, Finance, Newspaper, Manual, and Academia. This collection encompasses a diverse array of documents from both existing datasets and public web resources. For Law and Finance, we randomly sample documents from CUAD [13] and FinanceBench [16]. Documents in CUAD have complex layouts, while FinanceBench provides documents with large and complex tables. Manual documents are expanded from [23] using ManualsLib  [^1] https://www.manualslib.com/ to incorporate a variety of multimodal elements. For Academia, we gather papers from Arxiv  [^2] https://arxiv.org, and for Textbook, we source content from LibreTexts  [^3] https://libretexts.org/ and public resources, with a particular focus on formulas. For Newspaper, we collect New York Times from pages  [^4] https://static01.nyt.com/images/ following [40], covering dates from 2024-05-01 to 2024-08-31. For each collected document, we provide ground truth structured data. Specifically, we begin with parsing all documents using Mathpix  [^5] https://mathpix.com/ for structural data extraction. We then ask expert-level annotators to revise the results, ensuring fidelity to the original structure and content of PDFs while mitigating any style deviations from Mapthix. Detailed descriptions can be found in Appendix Sec. II.
3.2 Question-answer pair generation
@@
To systematically assess the impact of OCR results on RAG performance, our Q&A generation approach revolves around the diversity of evidence sources and realistic question-answer types representativeness for retrieval-augmented generation. Specifically, we provide the ground truth structured data of each document page to GPT-4o for generating Q&A based on multimodal document elements, including plain text, tables, and formulas. Each Q&A consists of the following fields: one page of the original PDF document, evidence context from ground truth structured data that provides the answer to the question, type of evidence (plain text, table, and formula), and the question and answer which are both derived from this evidence context. In this way, these Q&As can serve as a testbed for evaluating OCR results on multimodal document elements. To ensure relevance to real-world RAG applications, we require that each question include at least one specific entity name, minimizing ambiguity. The prompt template for the Q&A generation is provided in Appendix Sec. I.
Quality Control. For each Q&A, we first instruct GPT-4o to answer the question without access to the ground truth evidence context. This step automatically filters out questions that are answerable without retrieval. Next, We manually review and eliminate questions with ambiguous references, such as "How many models are compared in Table 3?", as these are challenging for RAG systems to resolve without clear context information. Finally, we ensure that each Q&A pair’s type of evidence aligns with the actual source context content. This multi-step quality control process helps guarantee that the Q&A dataset meets the requirements of diverse evidence sources and practical RAG applications.

### v2

@@ Abstract
Retrieval-augmented Generation (RAG) enhances Large Language Models (LLMs) by integrating external knowledge to reduce hallucinations and incorporate up-to-date information without retraining. As an essential part of RAG, external knowledge bases are commonly built by extracting structured data from unstructured PDF documents using Optical Character Recognition (OCR). However, given the imperfect prediction of OCR and the inherent non-uniform representation of structured data, knowledge bases inevitably contain various OCR noises. In this paper, we introduce OHRBench, the first benchmark for understanding the cascading impact of OCR on RAG systems. OHRBench includes 8,561 carefully selected unstructured document images from seven real-world RAG application domains, along with 8,498 Q&A pairs derived from multimodal elements in documents, challenging existing OCR solutions used for RAG. To better understand OCR’s impact on RAG systems, we identify two primary types of OCR noise: Semantic Noise and Formatting Noise and apply perturbation to generate a set of structured data with varying degrees of each OCR noise. Using OHRBench, we first conduct a comprehensive evaluation of current OCR solutions and reveal that none is competent for constructing high-quality knowledge bases for RAG systems. We then systematically evaluate the impact of these two noise types and demonstrate the trend relationship between the degree of OCR noise and RAG performance. Our OHRBench, including PDF documents, Q&As, and the ground truth structured data will be released to foster the development of OCR tailored to RAG and RAG systems that are resilient to OCR noise.
footnotetext: These authors contributed equally to this work.footnotetext: Corresponding author (heconghui@pjlab.org.cn).
@@ 1 Introduction
As an essential component of RAG systems, the knowledge base defines the scope and quality of documents that RAG can access. Given that a vast amount of real-world knowledge resides in unstructured documents, such as scanned PDFs, constructing an external knowledge base often relies on Optical Character Recognition (OCR)  [^1] We employ the General OCR concept for document parsing from GOT-OCR2.0 [43], which includes, text recognition, multimodal data extraction (table, formula, and chart recognition), and reading order restoration. to parse structured data from these unstructured PDF documents [48, 18]. For instance, MinerU [41] takes raw PDFs as input and extracts plain text, formulas, and tables into structured formats for subsequent RAG applications. However, imperfect predictions of OCR and non-uniform representations of parsing results impair the construction of a high-quality knowledge base for RAG. To be specific, despite advancements in OCR [41, 43, 3], even the leading model cannot achieve perfect accuracy across all scenarios [28, 48]. Furthermore, structural data like table can inherently be parsed in different representation, such as Markdown or LaTeX. These issues introduce OCR noise in parsing results and diminish the quality of the knowledge base. Considering RAG is sensitive to input noise [13, 45, 9], recent works race on downstream RAG components, including more precise retrievers [5, 30, 24] and more advanced LLMs [1, 11, 46, 13]. However, the quality of OCR-based external knowledge bases and its cascading impact on these downstream RAG components have received less attention, which highlights a critical but unaddressed gap: the absence of benchmarks to assess OCR’s cascading impact on each component and entire system of RAG.
Figure 1: Our OHRBench comprises documents from 7 domains, 9 challenging attributes for OCR, 4 types of Q&A tasks, and 5 Q&A evidence sources. Each number indicates the count of PDF pages with that attribute. Criteria for these attributes can be found in Appendix Sec. II.3
@@
Existing benchmarks either evaluate RAG holistically without fine-grained assessment [49], consider limited OCR solutions without accounting for the noise they introduce [18, 14]. Additionally, they lack documents that present more diverse OCR challenges, such as scanned historical, multilingual, and handwritten documents. To fill this gap, we introduce OHRBench, a question-answering benchmark designed to evaluate OCR’s cascading impact on each component and entire systems of RAG in two ways. First, we construct a document-based RAG Q&A dataset comprising complex, unstructured PDF documents from 7 RAG real-world application areas: Textbook, Law, Finance, Newspaper, Manual, Academic and Administration. As detailed in Tab. 1 and Fig. 1, we have collected 8,561 document images featuring attributes that challenges the creation of high-quality knowledge bases for RAG systems. We also provide diverse Q&A pairs which not only span realistic RAG tasks, including understanding, reasoning, and multi-page questions, but also features evidence sourced from key components of OCR in document parsing, making them ideal for assessing the OCR’s impact on RAG performance. Second, we identify two primary OCR noise types: Semantic Noise, resulting from prediction errors, and Formatting Noise, arising from diverse document element representation. By systematically introducing these noise types into documents, we generate perturbed structured data with varying degrees of noise, enabling further exploration of the quantitative relationship between OCR noise and RAG performance.
@@
With OHRBench, we first conduct a comprehensive benchmark on current OCR solutions, including pipeline-based OCR systems [41, 32], end-to-end OCR models [3, 43] and Vision-Language Models (VLMs) for OCR [2, 42, 7, 6]. We reveal that even the best OCR solutions exhibit a performance gap of 14% at least, compared to the ground truth structured data, facilitating the importance of mitigating OCR noise in RAG systems. Further experiments on different types of OCR noise uncover that Semantic Noise consistently exert a significant impact, while Formatting Noise affects specific retrievers and LLMs differently, offering valuable insights for developing RAG-tailored OCR solutions and noise robust models.
@@
  • We present OHRBench, a question-answering benchmark designed to evaluate the impact of OCR on RAG systems. OHRBench includes various unstructured PDF documents from seven RAG domains with ground truth structure data annotations and Q&A pairs spanning multiple RAG tasks with diverse source of evidences, posing challenges to the employment of current OCR solutions in RAG systems.
  • We conduct a comprehensive evaluation of current OCR solutions and reveal that none of them is competent for constructing high-quality knowledge bases for RAG systems.
  • We identify two primary types of OCR noise, including Semantic Noise and Formatting Noise, generate perturbed data with varying levels of noise and explore the trend relationship between the degree of OCR noise and RAG performance.
@@ 2.1 Retrieval-Augmented Generation
Retrieval-Augmented Generation (RAG) [23, 21, 34] integrates external knowledge into large language models (LLMs) to mitigate hallucinations. Although RAG technology enhances the generation capabilities of LLMs, it is notably sensitive to input noise. InfoRAG [45] characterizes this noise in RAG as incorrect and irrelevant content within retrieved text and reveals its impact on RAG performance. RAAT [13] further expands noise into relevant noise, counterfactual noise, and irrelevant types. However, these studies focus solely on chunk-level noise introduced during the retrieval stage and its effect on the generation capabilities of LLMs, leaving the impact of noise derived from OCR results unexplored. GARAG [9] examines typographical errors, a form of OCR noise, but its scope is limited to plain text using only synthetic data, overlooking the variety of OCR noise encountered in real-world RAG applications. In this paper, we reveal the impact of noise introduced during the OCR stage, offering a comprehensive analysis of its impact on RAG systems.
@@ 2.2 Document parsing with OCR
OCR-based document parsing is a promising solution for structured data extraction from unstructured documents, facilitating applications like RAG. Current OCR solutions can be summarized into three categories, pipeline-based systems [41, 32], end-to-end models [43, 3, 27], and employing VLMs for OCR [7, 42, 19] Pipeline-based systems decompose OCR into multiple subtasks, such as layout detection, text, formula, and table recognition, enabling fine-grained data extraction. End-to-end models take document images as input and output the overall recognition result in an end-to-end manner. Due to the achievement of VLMs on visual understanding, recent works have explored its application in OCR [28]. In this paper, we evaluate these OCR paradigms, examining their suitability for RAG applications across diverse, real-world document domains.
@@ 2.3 Benchmark and evaluation of Retrieval-Augmented Generation
Frameworks like RAGAS [12] and ARES [36] propose evaluating RAG systems based on context relevance, answer faithfulness, and answer relevance, using LLMs or fine-tuned discriminators for measurement. RGB [4] assesses the noise robustness, negative rejection, information integration, and counterfactual robustness of RAG in news data. MultiHop-RAG [37] focuses on multi-hop reasoning capabilities, while ClashEval [44] explores the context preference in conflicting evidence scenarios. However, these evaluations target specific components of RAG systems, and none of them discusses the impact of external knowledge base construction on RAG systems. Although UDA [18], VisRAG [47] and M3DocRAG [8] explore RAG’s effectiveness in document understanding, they consider limited OCR solutions, ignoring challenging documents and lacking analysis of different OCR noise types. In this paper, we introduce OHRBench to comprehensively investigate OCR noise’s impact on RAG systems.
@@
Documents $1,261$
- Domain $7$
- Total Pages $8,561$
- Avg.Tokens $1,034/\text{page}$
- Avg.Data Type $1.9/\text{page}$
Questions $8,498$
Avg.Question Token $17.9\pm 8.7$
Avg.Answer Token $3.5\pm 4.1$
@@
- Text $3,528\ (41.5\%)$
- Table $2,364\ (27.8\%)$
- Formula $1,267\ (14.9\%)$
- Chart $768\ (9.0\%)$
- Reading Order $691\ (8.1\%)$
@@
- String $4,171\ (54.4\%)$
- Numeric $3,004\ (35.3\%)$
- Yes/No $594\ (7.0\%)$
- List $483\ (6.9\%)$
(Task Type)
- Understanding $6,114\ (71.9\%)$
- Reasoning $2,384\ (28.1\%)$
- Single-page $7,656\ (90.1\%)$
- Multi-page $842\ (9.9\%)$
@@
Figure 2: Construction of OHRBench and evaluation protocol. (1) Benchmark Dataset: documents from seven domains, human-verified ground truth structured data, and Q&As from multimodal document elements. (2) RAG Knowledge Base: Current OCR results for benchmarking and perturbed data for assessment. (3) Evaluation of OCR impact on each component and the overall RAG system.
@@ 3 OHRBench
Our OHRBench consists of 1) a number of unstructured PDF documents from seven real-world RAG applications, Q&A pairs derived from multimodal document elements and ground truth structure data annotation, and 2) perturbed structured data based on ground truth with varying degrees of OCR noises.  Fig. 2 illustrates the construction of OHRBench. We will now delve into the details of each component.
@@ 3.1 Data collection
According to [48, 25], extracting structured data from multimodal document elements like formulas and tables poses significant challenges to current OCR solutions. Considering the practical application scenarios of RAG and the challenging field of OCR, we compile a PDF document collection representing seven common RAG application scenarios: Textbook, Law, Finance, Newspaper, Manual, Academic and Administration. This collection includes a diverse array of documents from both existing datasets and public web resources. Specifically, we first collect PDF documents from a wide range of existing datasets, including DUDE [40], OmniDocBench [31], FinanceBench [20], CUAD [16], and GNHK [22]. This results in a highly diverse PDF dataset that encompasses complex structured data and layouts, high text density, handwritten, scanned, and historical documents, as well as multilingual content (Chinese and English), which covers most challenges faced by OCR in document parsing. In addition to existing datasets, we further supplement our collection with documents from public resources to balance the distribution. We filter out the corrupted or license-restricted documents and finally curate a document dataset comprising 1,261 PDFs and 8,561 images. For each collected document, we manually categorize them into 7 domains and provide ground truth structured data. Specifically, we begin with parsing all documents using Mathpix [^2] https://mathpix.com/ for structural data extraction. We then ask expert-level annotators to revise the results, ensuring fidelity to the original structure and content of PDFs while mitigating any style deviations from Mapthix. Detailed descriptions of our selection and processing pipeline can be found in the Appendix  Sec. II.
3.2 Q&A pairs generation
@@
The process of extracting structured data from documents involves three key tasks: recognizing plain text; extracting multimodal document elements, including tables, formulas, and charts; and restoring reading order which includes multi-column and truncated paragraph merging. To systematically assess the impact of OCR results on RAG performance, our Q&A generation approach revolves around these 5 evidence sources and various realistic Q&A tasks. Specifically, we provide the ground truth structured data of each document page to GPT-4o and prompt it to generate Q&A based on important components in document parsing, including plain text, tables, formulas, and charts. For questions related to reading order, we identify paragraphs that require merging and instruct GPT-4o to create questions that necessitate combining these paragraphs for a complete answer. We generate both understanding questions, which only require extracting specific information, and reasoning questions, which involve arithmetic operations, comparisons, or synthesizing information across multiple sections. For multi-page Q&A, we derive them from both single-page Q&A and ground truth structured data from different pages that share the same entity name, recognized with spaCy [17]. Detailed process and the prompt template for the Q&A generation are provided in Appendix Sec. I. Each Q&A consists of the following fields: one page of the original PDF document, evidence context from ground truth structured data that provides the answer to the question, type of evidence (plain text, table, formula, chart and reading order), and the question and answer which are both derived from this evidence context. In this way, these Q&As can serve as a testbed for evaluating OCR results on multimodal document elements.
Quality Control. The quality of Q&A pairs generated by a large language model (LLM) can vary significantly. To address this issue, we apply three data selection criteria to ensure high-quality Q&As: (1) compatibility with realistic RAG applications, (2) faithfulness to task definitions and (3) correctness. We incorporate both heuristic methods and prompting LLM for auto data filtering:
  • Compatibility to RAG Applications. Questions should be context-independent and not answerable by the model’s internal knowledge. We collect keywords from existing context-dependent questions, such as “according to the document”, as heuristic rules. Following [47], we also employ LLMs to classify context dependence for further filtering. Questions answerable without retrieval are excluded by instructing LLMs to answer without access to the evidence context.
  • Faithfulness to Task Definition. We ensure questions align with their task definition (reasoning or understanding) using LLMs to judge and that evidence sources match the context using heuristic rules. In addition, for multi-page question, we provide single evidence context and use LLMs’ responses to filter out answerable questions.
  • Correctness. We verify the accuracy of both evidence context and answers. In specific, we provide LLMs with oracle evidence contexts and sample answers repeatedly to filter Q&A below a certain correctness threshold.
Finally, we manually check against the criteria to ensure the quality. The LLMs used in our verification include GPT-4o and DeepSeek-V3 [26], where we find that DeepSeek-V3 achieve similar performance compared to GPT-4o on Q&A verification. A detailed description can be found in the Appendix Sec. II.5. This multi-step quality control ensures the Q&A dataset meets diverse evidence source requirements and practical RAG applications. Ultimately, we filter out 8498 high-quality Q&As from 15317 candidates.